OpenStudy (anonymous):

Systems of Equations in 5 Variables

OpenStudy (anonymous):

2a-3b+c-d-2e=-7 a+2b+4c+d+5e=11 2a+3b+3c-4d+3e=7 a+b-4c+2d+e=6 5a+3b-8c+3d+4e=18

OpenStudy (zzr0ck3r):

Gaussian elimination...

OpenStudy (anonymous):

Or I could just use my calculator..

OpenStudy (zzr0ck3r):

lol, nah you should row reduce that

OpenStudy (zzr0ck3r):

looks like a blast

OpenStudy (anonymous):

Good thing that's not MY homework

OpenStudy (experimentx):

the equation is linear ... it's not that bad as you expect. Use Cramers rule or Gaussian elimination

OpenStudy (zzr0ck3r):

why is this still open?

OpenStudy (zzr0ck3r):

@FutureMathProfessor

OpenStudy (anonymous):

Because no one has done it yet

OpenStudy (zzr0ck3r):

I told you how and you said "or you could do it in your calculator" so I told you how to do it, and you said you don't want to do it. so close the question, we are not here to give answers to questions, we are here to help. If this is not helping the close it.....